Fermi National Laboratory
 

Accelerator Update for October 19, 2007 - October 22, 2007

 

Friday October 19

The day shift began with the Antiproton Source (Pbar) stacking (100E10), with Operations monitoring a Recycler Stash (30E10), with the Main Injector (MI), Recycler, Pbar, and Tevatron (TeV) conducting studies, with the SciBooNE / MiniBooNE experiments taking beam, with the Neutron Therapy Facility (NTF) treating patients, and with the TeV in a dry squeeze.

At 8:41 AM, Operations reported that Pbar had stopped stacking after having 1.5 hours of problems. Operators made some adjustments and stacking resumed about 20 minutes later.

A Sci/MiniBooNE expert reported at 9:09 AM that a beam monitor (VP869) was giving bad readbacks. He said he would take care of this.

At 9:48 AM, a TeV kicker (PAK3) prefired and aborted the beam. Operators put the TeV into a dry squeeze while experts investigated. Much later an expert reported that it was a power supply problem.

Recycler experts began tuning their beamline at 10:39 AM in preparation for transferring antiprotons from Pbar.

NTF completed treating their last patient for the day at 11:22 AM.

Operations reported at 12:11 PM that 15 transfers of antiprotons from Pbar to the Recycler were attempted and most completed. Pbar resumed stacking at 12:48 PM.

At 1:04 PM, MI-30 suffered a vacuum burst.

At 1:27 PM, a TeV expert reported that the PAK3 power supply repair would take about four hours.

TeV experts resumed studies around 2:10 PM.

At 2:34 PM, Operations reported a TECAR fast bypass at TeV sector A3. There was no quench. Later, experts reported that they found a problem with a TeV D2 power supply connector.

MI experts began aperture scans at 3:11 PM.

The TeV resumed ramping at 4:11 PM.

Recycler/Pbar experts made the first of two antiproton transfers at 6 PM. The Pbar emittances were big. Experts stopped the transfers and investigated. Later they reported that the problem was due to the Pbar horizontal dampers.

At 8 PM, Operations took the TeV separators out of pulsing mode for four hours per instructions.

D0 began an access at 9:10 PM.

Operations made the first of two Recycler stashes at 10:06 PM. They reported that there was still an emittance blow up problem.

D0 came out of access at 11:54 PM.

Saturday October 20

The midnight shift began with Pbar stacking (26.9E10), with Operations monitoring the stash (72.5E10), with TeV experts conducting studies, and with Sci/MiniBooNE taking beam.

At 1:22 AM, the TeV suffered a quench at D11 while conducting a low beta magnet squeeze.

Operations began a Recycler stash at 2:25 AM. After the stash, Pbar Operations halted stacking and put the machine into standby.

At 4:36 AM, the Cryo Coordinator reported a helium leak from a wet engine at the TeV sector F3 cryo system building. Experts were called in to fix the leak.

Operations turned on the Recycler’s Electron Cooling at 8:11 AM per expert’s request.

Cryo system experts reported at 9:11 AM that the TeV sector F3 wet engine was operating and cooldown had begun.

Electron Cooling experts began a study period at 9:54 PM.

Operations put the TeV through a dry squeeze at 11:38 AM.

At 2:34 PM, Operations reported that a Booster RF station (BRF16) wasn’t operating properly. Operators bypassed the station. Later, an expert reported that the station was sparking and an access would be needed to repair.

At 2:38 PM, a MI RF station (MIRF4) wasn’t operating properly. Operators bypassed the station.

Pbar experts began a stacking study at 4:25 PM.

At 7:43 PM, Operations reported problems with Pbar core stability. Experts investigated.

At 10:55 PM, Operations reported that the P2 beam permit had dropped due to a power supply (I:HTF17) trip. Operators replaced the power supply’s regulator, but that didn’t resolve the problem. Operators continued to investigate.

Sunday October 21

The midnight shift began with a Pbar expert stacking and conducting studies, with Operations monitoring the stash, with TeV and MI experts conducting studies, and with Sci/MiniBooNE taking beam.

Operators reported at 12:27 AM that they reseated the DC power input to the HTF17 regulator and it began working.

Operations began stashing antiprotons at 3:13 AM.

At 4:44 AM, BRF19 tripped off for the first of six times in about 45 minutes. Operators reset it. An expert came in at 7:56 AM and reported that a new gate delay module had been installed. He switched it back to the old gate.

At 3:22 PM, Operations reported a power glitch. The Booster West Anode Power Supply tripped off as well has several Linac devices. The Duty Electrician reported that he couldn’t find any indication of a problem at the Master Substation. Everything reset.

At 9:15 PM, the TeV suffered a quench at sector A11 due to TRF7 tripping off with TRF1 already off.

At 10:19 PM, Operations reported that the chromaticity studies had caused beam to accumulate in the abort gap, then TRF7 tripped and since TRF1 was already off, the abort kickers fired. Since there was a lot of beam in the gap it was a sloppy abort and that caused the low beta magnets to quench.

Monday October 22

The midnight shift began with a Pbar expert stacking, with Operations monitoring the stash, with TeV and MI experts conducting studies, and with Sci/MiniBooNE taking beam.

TeV experts put the TeV through the first of three high-energy ramps at 1:03 AM.

Operations put the TeV through a dry squeeze at 1:21 AM.

Operations began stashing the Recycler at 1:46 AM.

Operations reported at 2:12 AM that TeV experts had parsed the squeeze on the central orbit and used the Tevatron Electron Lens (TEL1) to scrub the abort gap.

Operations stashed the Recycler at 7:02 AM.

The Plans for Monday

The current status for today is as follows: NTF is treating patients; the Linac, Booster, and MI are all operating normally, the Recycler has a stash (235E10), Pbar is stacking (21mA), NuMI is still in shutdown mode, Switchyard 120 would like to begin power supply testing later this week, and the Sci/MiniBooNE experiments are taking beam. There will be a 12 to 16 hour shutdown for D0 tomorrow (10/23/07) to open their detector and make repairs. Many of the machines will use this opportunity to access tunnels and also make repairs.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/22/2007    email Fermilab